OpenStudy (zyberg):

Prove that if numbers a and b are coprime, then there is such natural number n, that a^n =(modulus equal sign)= 1 (mod b)

OpenStudy (zyberg):

As I understand, I need to try looking at a/b and then a^n/b. What gets me really confused is how would I predict reminders of a^n/b? Or is it not needed?

OpenStudy (welshfella):

a^n = 1 (mod b) (modular equals) as a consequence of this i know we can say the following a^n - 1 = bk for some integer k but what next i dont know.

OpenStudy (loser66):

\(a^n\equiv 1\) mod b , is what we have to prove. @welshfella

OpenStudy (welshfella):

yeah of course I just wondered if that transformation might help in the proof

OpenStudy (mathmate):

@Zyberg FYI \(\equiv\) is written in LaTeX as \equiv .

OpenStudy (sparrow2):

i remember that somewhere i learnt this: